# Parameters Optimization in Trajectory Planning Using Diffrentiable Convex Programing

### v1

@@ Abstract
Sequential convex programming has been established as an effective framework for solving nonconvex trajectory planning problems. However, its performance is highly sensitive to problem parameters, including trajectory variables, algorithmic hyperparameters, and physical vehicle parameters. This paper introduces a differentiable sequential convex programming framework that integrates differentiable convex optimization with SCP to enable end-to-end parameter optimization. By deriving first-order sensitivity relations of second-order cone programming solutions with respect to problem data, exact gradients of trajectory performance metrics with respect to arbitrary parameters are obtained and propagated through SCP iterations. The effectiveness of the proposed framework is validated through three representative applications: optimal terminal-time prediction for powered landing, trust-region penalty optimization in SCP subproblems, and surface-to-mass ratio optimization for hypersonic gliding vehicles. Simulation results show that the proposed framework enables reliable gradient-based parameter learning and significantly improves numerical performance, convergence behavior, and design efficiency. These results indicate that differentiable sequential convex programming framework provides a powerful and general tool for vehicle design, mission optimization, and hyperparameter selection in aerospace trajectory planning.
@@ 1 Introduction
Trajectory planning plays a central role in the performance and success of aerospace missions. Its computational efficiency and convergence reliability directly determine whether guidance and control can be executed in real time [1, 2]. Sequential convex programming (SCP) is an effective approach for trajectory optimization in applications such as powered landing and hypersonic vehicle entry[3, 4]. Recent advances in SCP have improved convergence and runtime through enhanced constraint handling, relaxation techniques[5], and discretization strategies[6]. Beyond algorithmic improvements, however, a wide range of problem parameters, such as initial trajectory guesses, trust-region penalty coefficients, normalization scales, and aerodynamic shape parameters, exert substantial influence on the efficiency and success rate of SCP-based trajectory planning. The systematic and quantitative analysis of how these parameters affect convergence behavior and mission performance remains limited. Motivated by this gap, the present study integrates differentiable convex optimization with SCP to construct a differentiable sequential convex programming (DSCP) framework capable of computing end-to-end sensitivities from problem parameters to trajectory-level performance metrics, thereby enabling direct parameter optimization.
@@
To investigate parameter influence in a structured manner, the parameters involved in SCP-based trajectory planning are categorized into three classes. The first class consists of the optimization variables intrinsic to the trajectory planning problem. To reduce problem dimension or enable real-time approximation, several works employ neural networks to learn subsets of trajectory variables. For instance, predicting terminal time for powered landing from SCP-generated datasets [7], generating initial guess trajectories for SCP [8], or approximating key quantities derived from optimality conditions [9]. Reinforcement-learning approaches can also learn control strategies [10, 11], but these methods focus on feedback decision-making rather than producing full trajectory profiles.
@@
The second category includes algorithmic parameters used in constructing SCP subproblems, such as trust-region penalties, normalization coefficients, and constraint penalty parameters, which are significant in the convergence behavior of SCP. In particular, appropriate normalization improves the conditioning of the Karush-Kuhn-Tucker (KKT) system [12, 13]. Meanwhile, trust-region mechanisms are essential to ensuring the validity of linearization and the feasibility of subproblems, while also influencing the convergence rate[2]. Existing studies have proposed various strategies, including dual-variable-based detection of trajectory oscillations for penalization [14], analysis of soft trust-region SCP showing that first-order penalties may not ensure optimal convergence and the design of higher-order penalties [15], and hybrid-order trust-region formulations that balance low-order speed and high-order optimality [16]. In [17], it is pointed out that selecting appropriate soft trust-region parameters remains a major design difficulty for nonlinear programming. To address this issue, [18] proposes an SCP formulation based on a high-dimensional merit function. By embedding the merit function directly into the optimization process, this method removes the need for explicit selection of soft trust-region parameters and thereby alleviates the associated tuning burden. In contrast, [19] introduced an alternative approach by separating nonconvex and convex constraints and developing an online trust-region adaptation scheme. This autotuned framework automatically adjusts trust-region bounds during the SCP iterations, improving convergence behavior and reducing reliance on manually chosen trust-region parameters.
@@
The third category consists of mission-level parameters in trajectory planning or physical parameters associated with the vehicle dynamics[20]. In [21, 22], a bilevel optimization framework was employed to determine the optimal switching point from the reentry phase to the powered-descent phase, treated as a key mission parameter. In the context of conceptual and preliminary vehicle design, aerodynamic-shape optimization typically requires extensive aerodynamic databases. Each aerodynamic model must be evaluated through comprehensive trajectory simulations and control-performance assessments. To reduce the computational cost of Computational Fluid Dynamics (CFD), existing acceleration techniques rely heavily on surrogate modeling for aerodynamic prediction. For example, [23] used transfer learning and model fine-tuning to learn aerodynamic characteristics from CFD data, thereby accelerating flight-trajectory simulation. Furthermore, [24] integrated surrogate models with multiobjective optimization to design a de-spin actuator, simultaneously addressing requirements on flight stability, range capability, and other key performance metrics. Despite these advances, there remains a lack of methods that directly leverage trajectory planning simulations to optimize aerodynamic parameters themselves. In particular, efficient approaches that treat trajectory-performance metrics as objective functions for aerodynamic-parameter tuning are still sparsely explored.
@@
To address the challenges associated with predicting and designing the three categories of parameters described above, the differentiable convex optimization approach adopted in this work enables the computation of exact gradients of trajectory performance metrics with respect to these parameters. This makes gradient-based optimization applicable to all three classes of parameters. Existing research in differentiable convex optimization generally derives the Jacobian of the optimal solution with respect to problem data through the KKT conditions or primal-dual residual systems. OptNet formulated a differentiable structure for quadratic programs (QPs) and developed a parallelizable QP solver that supports efficient batched forward and backward passes[25]. DiffCP[26] and DiffQCP[27] extended this idea to general cone programs by solving coupled primal-dual residual equations to obtain parameter sensitivities. These methods provide a unified treatment for differentiable conic optimization. BPQP further exploited the KKT structure and the characterization of active sets, reformulating the backward pass as another QP [28]. This construction substantially improves the computational efficiency of the gradient propagation. Based on these differentiable optimization layers, prior work has integrated differentiable model predictive control into reinforcement-learning frameworks [29]. This integration leverages the exploration capabilities of reinforcement learning together with the differentiability of optimization, improving policy-learning performance. Differentiable convex optimization has also been applied in trajectory planning of unmanned aerial vehicle (UAV)[30]. In this line of work, a single differentiable optimization layer is embedded within a neural network to train a trajectory planning model, enabling real-time UAV trajectory generation from depth maps.
@@ 2 Problem Statement
In this section, we present the trajectory planning framework based on SCP, including the discretization scheme, the linearization procedure, and the resulting SOCP formulation. We further clarify the categories of parameters involved in the SCP process. Consider the trajectory planning problem with time-invariant system dynamics
@@
where $\boldsymbol{x}$, $\boldsymbol{u}$ and $J$ denote the state and control variables and the performance index. To transform the continuous optimal control problem into a finite-dimensional optimization problem, all the constraints must be discretized, especially the system dynamics equation. In this work, we adopt a trapezoidal-rule-inspired discretization method, following our work in [31]. Let $N$ denote the number of discrete time intervals, each with duration $T$. The discrete form of the dynamics constraints is given by
$\boldsymbol{x}[n+1]-\boldsymbol{x}[n]=T\boldsymbol{f}\left(\bar{\boldsymbol{x}},\bar{\boldsymbol{u}}\right),n\in\left[0,N-1\right]$ (2)
@@
$\bar{\boldsymbol{x}}=\frac{\boldsymbol{x}[n+1]+\boldsymbol{x}[n]}{2},\bar{\boldsymbol{u}}=\frac{\boldsymbol{u}[n+1]+\boldsymbol{u}[n]}{2}$ (3)
@@
$\delta\boldsymbol{x}[n+1]-\delta\boldsymbol{x}[n]+\boldsymbol{x}_{\text{ref}}[n+1]-\boldsymbol{x}_{\text{ref}}[n]=T_{\text{ref}}\boldsymbol{f}_{\text{ref}}+T_{\text{ref}}\boldsymbol{A}_{\text{ref}}\delta\bar{\boldsymbol{x}}[n]+T_{\text{ref}}\boldsymbol{B}_{\text{ref}}\delta\bar{\boldsymbol{u}}[n]+\delta T\boldsymbol{f}_{\text{ref}}$ (4)
@@
$\boldsymbol{A}_{\text{ref}}=\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\bar{\boldsymbol{x}}_{\text{ref}},\bar{\boldsymbol{u}}_{\text{ref}}},\boldsymbol{B}_{\text{ref}}=\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\bar{\boldsymbol{x}}_{\text{ref}},\bar{\boldsymbol{u}}_{\text{ref}}},\boldsymbol{f}_{\text{ref}}=\boldsymbol{f}(\bar{\boldsymbol{x}}_{\text{ref}},\bar{\boldsymbol{u}}_{\text{ref}})$ (5)
@@
$\boldsymbol{r}_{\text{ref}}$ $=\left(\frac{T_{\text{ref}}\boldsymbol{A}_{\text{ref}}}{2}+\boldsymbol{I}\right)\delta\boldsymbol{x}[n]+\left(\frac{T_{\text{ref}}\boldsymbol{A}_{\text{ref}}}{2}-\boldsymbol{I}\right)\delta\boldsymbol{x}[n+1]$ (6)
$+\frac{T_{\text{ref}}\boldsymbol{B}_{\text{ref}}}{2}\delta\boldsymbol{u}[n]+\frac{T_{\text{ref}}\boldsymbol{B}_{\text{ref}}}{2}\delta\boldsymbol{u}[n+1]+\boldsymbol{f}_{\text{ref}}\delta T$
where $\boldsymbol{r}_{\text{ref}}=(\boldsymbol{x}_{\text{ref}}[n+1]-\boldsymbol{x}_{\text{ref}}[n])-T_{\text{ref}}\boldsymbol{f}_{\text{ref}}$ denotes the discretization residual associated with the reference trajectory. This term mitigates the adverse effects that may arise in subsequent iterations when the reference trajectory does not fully satisfy the system dynamics. Using the same procedure as for the dynamics constraints, the cost function and path constraints in Eq.(1) are discretized and linearized into the following form:
@@
$N_{\mathrm{SOC}}$ denotes the number of second-order cone constraints, and $m_{i}$ represents the dimension of the $i$-th cone. The vector $\boldsymbol{\theta}$ aggregates all parameters involved in the SCP subproblem, including the reference trajectory $\boldsymbol{z}_{\text{ref}}$, the SCP-specific parameters $\boldsymbol{\theta}_{\text{SCP}}$ and the trajectory planning parameters $\boldsymbol{\theta}_{\text{TP}}$. By repeatedly solving the subproblem in Eq.(12) and updating the reference trajectory, SCP ultimately converges to the optimal solution of the problem in Eq.(1). Although SCP is widely recognized as an effective method for trajectory planning, its convergence efficiency is highly sensitive to parameter selection, which has attracted increasing research attention in recent years. In addition, fixing part of the nonconvex decision variables prior to optimization, thereby treating them as parameters during the solution process, has proven to be an effective means of accelerating computation. Motivated by these observations, this work investigates the influence of parameter optimization on the cost function of SCP-based trajectory planning. A differentiable convex optimization framework is integrated into the SCP scheme, enabling end-to-end learning of parameters through a differentiable SCP pipeline.
3 Differentiable Sequencial Convex Programing

### v2

# Parameter Optimization in Trajectory Planning via Differentiable Convex Programming

@@ Abstract
Sequential convex programming has been established as an effective framework for solving nonconvex trajectory planning problems. However, its performance is highly sensitive to problem parameters, including trajectory variables, algorithmic hyperparameters, and physical vehicle parameters. This paper introduces a differentiable sequential convex programming framework that integrates differentiable convex optimization with sequential convex programming to enable end-to-end parameter optimization. By deriving first-order sensitivity relations of second-order cone programming solutions with respect to problem data, exact gradients of trajectory performance metrics with respect to arbitrary parameters are obtained and propagated through iterations. The effectiveness of the proposed framework is validated through three representative applications: optimal terminal-time prediction for powered landing, trust-region penalty optimization in subproblems, and surface-to-mass ratio optimization for hypersonic gliding vehicles. Simulation results show that the proposed framework enables reliable gradient-based parameter learning and significantly improves numerical performance, convergence behavior, and design efficiency. These results indicate that the differentiable sequential convex programming framework provides a powerful and general tool for vehicle design, mission optimization, and hyperparameter selection in aerospace trajectory planning.
@@ 1 Introduction
Trajectory planning plays a central role in the performance and success of aerospace missions. Its computational efficiency and convergence reliability directly determine whether guidance and control can be executed in real time [heReviewDatadrivenComputational2025, wangSurveyConvexOptimization2024]. Sequential convex programming (SCP) is an effective approach for trajectory optimization in applications such as powered landing and hypersonic vehicle entry [lu2025model, liu2016entry]. Recent advances in SCP have improved convergence and runtime through enhanced constraint handling, relaxation techniques [szmukSuccessiveConvexificationFuelOptimal2016], and discretization strategies [kimOptimalMidcourseGuidance2023a]. Beyond algorithmic improvements, however, a wide range of problem parameters, such as initial trajectory guesses, trust-region penalty coefficients, normalization scales, and aerodynamic shape parameters, exert substantial influence on the efficiency and success rate of SCP-based trajectory planning. The systematic and quantitative analysis of how these parameters affect convergence behavior and mission performance remains limited. Motivated by this gap, the present study integrates differentiable convex optimization with SCP to construct a differentiable sequential convex programming (DSCP) framework capable of computing end-to-end sensitivities from problem parameters to trajectory-level performance metrics, thereby enabling direct parameter optimization.
@@
To investigate parameter influence in a structured manner, the parameters involved in SCP-based trajectory planning are categorized into three classes. The first class consists of the optimization variables intrinsic to the trajectory planning problem. To reduce problem dimension or enable real-time approximation, several works employ neural networks (NNs) to learn subsets of optimization variables. For example, NNs are trained to predict terminal time for powered landing from datasets generated by lossless convex programming [liFreeFinalTimeFuelOptimal2022a]. In [youLearningBasedOnboardGuidance2021a, YOU2022129], key parameters of the powered landing problem are derived from the first-order necessary conditions of optimal control, and neural networks are used to approximate these parameters, which significantly reduces the computational burden of online trajectory planning. Reinforcement-learning approaches can also learn control strategies [takahashiAutonomousReconnaissanceTrajectory2023, yangReinforcementLearningBasedRobustGuidance2024a], but these methods focus on feedback decision-making rather than producing full trajectory profiles.
@@
The second category includes algorithmic parameters used in constructing SCP subproblems, such as trust-region penalties, normalization coefficients, and constraint penalty parameters, which are significant in the convergence behavior of SCP. In particular, appropriate normalization improves the conditioning of the Karush-Kuhn-Tucker (KKT) system [domahidiECOSSOCPSolver2013, kamathOptimalPreconditioningOnline2025]. Meanwhile, trust-region mechanisms are essential to ensuring the validity of linearization and the feasibility of subproblems, while also influencing the convergence rate [wangSurveyConvexOptimization2024]. Existing studies have proposed various strategies, including dual-variable-based detection of trajectory oscillations for penalization [baeNewTrustRegionConstraint2025], analysis of soft trust-region SCP showing that first-order penalties may not ensure optimal convergence and the design of higher-order penalties [xieHigherOrderSoftTrustRegionBasedSequential2023], and hybrid-order trust-region formulations that balance low-order speed and high-order optimality [xieHybridorderSoftTrust2024]. In [fletcherNonlinearProgrammingPenalty2002], it is pointed out that selecting appropriate soft trust-region parameters remains a major design difficulty for nonlinear programming. To address this issue, [xieHighDimensionalMeritFunctionBasedSequential2025] proposes an SCP formulation based on a high-dimensional merit function. By embedding the merit function directly into the optimization process, this method removes the need for explicit selection of soft trust-region parameters and thereby alleviates the associated tuning burden. In contrast, [mceowenAutotunedPrimalDual2025] introduces an alternative approach by separating nonconvex and convex constraints and developing an online trust-region adaptation scheme. This autotuned framework automatically adjusts trust-region bounds during the SCP iterations, improving convergence behavior and reducing reliance on manually chosen trust-region parameters.
@@
The third category consists of mission-level parameters in trajectory planning or physical parameters associated with the vehicle dynamics [hoburgGeometricProgrammingAircraft2012, zheng2025model]. In [luPropellantOptimalPoweredDescent2023, luFastRobustOptimization2024], a bilevel optimization framework is employed to determine the optimal switching point from the reentry phase to the powered-descent phase, treated as a key mission parameter. In the context of conceptual and preliminary vehicle design, aerodynamic-shape optimization typically requires extensive aerodynamic databases. Each aerodynamic model must be evaluated through comprehensive trajectory simulations and control-performance assessments. To reduce the computational cost of Computational Fluid Dynamics (CFD), existing acceleration techniques rely heavily on surrogate modeling for aerodynamic prediction. For example, [jiUnsteadyAerodynamicModeling2024] used transfer learning and model fine-tuning to learn aerodynamic characteristics from CFD data, thereby accelerating flight-trajectory simulation. Furthermore, [luoIntegratedOptimizationDespin2025] integrated surrogate models with multiobjective optimization to design a de-spin actuator, simultaneously addressing requirements on flight stability, range capability, and other key performance metrics. Despite these advances, there remains a lack of methods that directly leverage trajectory planning simulations to optimize aerodynamic parameters themselves. In particular, efficient approaches that treat trajectory-performance metrics as objective functions for aerodynamic-parameter tuning are still sparsely explored.
@@
To address the challenges associated with predicting and designing the three categories of parameters described above, the differentiable convex optimization approach adopted in this work enables the computation of exact gradients of trajectory performance metrics with respect to these parameters. This makes gradient-based optimization applicable to all three classes of parameters. Existing research in differentiable convex optimization generally derives the Jacobian of the optimal solution with respect to problem data through the KKT conditions or primal-dual residual systems. OptNet formulated a differentiable structure for quadratic programs (QPs) and developed a parallelizable QP solver that supports efficient batched forward and backward passes [amosOptNetDifferentiableOptimization2021]. DiffCP [agrawalDifferentiableConvexOptimization2019] and DiffQCP [healeyDifferentiatingQuadraticCone2025] extended this idea to general cone programs by solving coupled primal-dual residual equations to obtain parameter sensitivities. These methods provide a unified treatment for differentiable conic optimization. BPQP further exploited the KKT structure and the characterization of active sets, reformulating the backward pass as another QP [panBPQPDifferentiableConvex]. This construction substantially improves the computational efficiency of the gradient propagation. Based on these differentiable optimization layers, prior work has integrated differentiable model predictive control into reinforcement-learning frameworks [romeroActorCriticModelPredictive2024]. This integration leverages the exploration capabilities of reinforcement learning together with the differentiability of optimization, improving policy-learning performance. Differentiable convex optimization has also been applied in trajectory planning of unmanned aerial vehicles (UAVs) [jiangSelfSupervisedLearningApproach2025]. In this line of work, a single differentiable optimization layer is embedded within a neural network to train a trajectory planning model, enabling real-time UAV trajectory generation from depth maps.
@@ 2 Problem Statement
In this section, the trajectory planning framework is presented based on SCP, including the discretization scheme, the linearization procedure, and the resulting SOCP formulation. We further clarify the categories of parameters involved in the SCP process. Consider the trajectory planning problem with time-invariant system dynamics
@@
where $\boldsymbol{x}$, $\boldsymbol{u}$ and $J$ denote the state and control variables and the performance index. To transform the continuous optimal control problem into a finite-dimensional optimization problem, all the constraints must be discretized, especially the system dynamics equation. In this work, a trapezoidal-rule-inspired discretization method is adopted, following our work in [zhangReentryTrajectoryOptimization2025]. Let $N$ denote the number of discrete time intervals, each with duration $T$. The discrete form of the dynamics constraints is given by
$\boldsymbol{x}[n+1]-\boldsymbol{x}[n]=T\boldsymbol{f}\left(\bar{\boldsymbol{x}}[n],\bar{\boldsymbol{u}}[n]\right),n\in\left[0,N-1\right]$ (2)
@@
$\bar{\boldsymbol{x}}[n]=\frac{\boldsymbol{x}[n+1]+\boldsymbol{x}[n]}{2},\bar{\boldsymbol{u}}[n]=\frac{\boldsymbol{u}[n+1]+\boldsymbol{u}[n]}{2}$ (3)
@@
$\delta\boldsymbol{x}[n+1]-\delta\boldsymbol{x}[n]+\boldsymbol{x}_{\text{ref}}[n+1]-\boldsymbol{x}_{\text{ref}}[n]=T_{\text{ref}}\boldsymbol{f}_{\text{ref}}+T_{\text{ref}}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\text{ref}}\delta\bar{\boldsymbol{x}}[n]+T_{\text{ref}}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\text{ref}}\delta\bar{\boldsymbol{u}}[n]+\delta T\boldsymbol{f}_{\text{ref}}$ (4)
@@
$\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\text{ref}}=\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\bar{\boldsymbol{x}}_{\text{ref}},\bar{\boldsymbol{u}}_{\text{ref}}},\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\text{ref}}=\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\bar{\boldsymbol{x}}_{\text{ref}},\bar{\boldsymbol{u}}_{\text{ref}}},\boldsymbol{f}_{\text{ref}}=\boldsymbol{f}(\bar{\boldsymbol{x}}_{\text{ref}},\bar{\boldsymbol{u}}_{\text{ref}})$ (5)
@@
$\boldsymbol{d}_{\text{ref}}$ $=\left(\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\text{ref}}+\boldsymbol{I}\right)\delta\boldsymbol{x}[n]+\left(\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\text{ref}}-\boldsymbol{I}\right)\delta\boldsymbol{x}[n+1]$ (6)
$+\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\text{ref}}\delta\boldsymbol{u}[n]+\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\text{ref}}\delta\boldsymbol{u}[n+1]+\boldsymbol{f}_{\text{ref}}\delta T$
where $\boldsymbol{d}_{\text{ref}}=(\boldsymbol{x}_{\text{ref}}[n+1]-\boldsymbol{x}_{\text{ref}}[n])-T_{\text{ref}}\boldsymbol{f}_{\text{ref}}$ denotes the discretization residual associated with the reference trajectory. This term mitigates the adverse effects that may arise in subsequent iterations when the reference trajectory does not fully satisfy the system dynamics. Using the same procedure as for the dynamics constraints, the cost function and path constraints in Eq.(1) are discretized and linearized into the following form:
@@
$N_{\mathrm{SOC}}$ denotes the number of second-order cone constraints, and $m_{i}$ represents the dimension of the $i$-th cone. The vector $\boldsymbol{\theta}$ aggregates all parameters involved in the SCP subproblem, including the reference trajectory $\boldsymbol{z}_{\text{ref}}$, the SCP-specific parameters $\boldsymbol{\theta}_{\text{SCP}}$ and the trajectory planning parameters $\boldsymbol{\theta}_{\text{TP}}$. By repeatedly solving the subproblem in Eq.(12) and updating the reference trajectory, SCP ultimately converges to the optimal solution of the problem in Eq.(1). Although SCP is widely recognized as an effective method for trajectory planning, its convergence efficiency is highly sensitive to parameter selection, which has attracted increasing research attention in recent years [xieHigherOrderSoftTrustRegionBasedSequential2023, xieHybridorderSoftTrust2024]. In addition, fixing part of the nonconvex decision variables prior to optimization, thereby treating them as parameters during the solution process, has proven to be an effective means of accelerating computation [liFreeFinalTimeFuelOptimal2022a, doi:10.2514/6.2024-1760]. Motivated by these observations, this work investigates the influence of parameter optimization on the SCP-based trajectory planning. A differentiable convex optimization framework is integrated into the SCP scheme, enabling end-to-end learning of parameters through a differentiable SCP pipeline.
3 Differentiable Sequential Convex Programming